# Enhancing Multimodal Medical Image Classification using Cross-Graph Modal Contrastive Learning

### v1

@@ Abstract
Neurodegenerative, SPECT, Contrastive learning, Multimodal fusion, Classification, Cross-graph modal graph learning.
@@ 6.4 Objective Function Ablation
We propose that a well-designed contrastive loss within the cross-graph modal objective function can effectively integrate feature spaces during the overall model optimization process. To evaluate CGMCL’s multimodal fusion performance, we tested various combinations of loss functions. These include the binary loss for single graph module fusion ($\mathcal{L}_{I}$ and $\mathcal{L}_{C}$), the addition of contrastive loss ($\mathcal{L}_{contrastive}$), and the comprehensive weighted loss CGMCL ($\mathcal{L}_{CGMCL}$). The experimental results in Fig. 4 show that $\mathcal{L}_{CGMCL}$ outperforms the combination of $\mathcal{L}_{I}$ + $\mathcal{L}_{C}$ by 1.2% in AUC and 2.4% in ACC for the MA vs. Abnormal classification task. Similarly, for the MA vs. Normal task, $\mathcal{L}_{CGMCL}$ shows improvements of 1.1% in both AUC and ACC. These results indicate that the similarity between the two modalities for these subtypes is nearing the model’s convergence range. More importantly, the ability to distinguish early-stage PD is critical. In the Normal vs. MA classification task, $\mathcal{L}_{CGMCL}$ surpasses the next best-performing unweighted model $\mathcal{L}_{CGMCL}$ by 7.1% in both AUC and ACC. This substantial improvement highlights CGMCL’s superior capability in capturing latent patterns characteristic of early-stage PD within the fusion space.

### v2

@@ Abstract
Neurodegenerative disease, SPECT, Contrastive learning, Multimodal fusion, Classification, Cross-graph modal graph learning.
@@ 6.4 Objective Function Ablation
We propose that a well-designed contrastive loss within the cross-graph modal objective function can effectively integrate feature spaces during the overall model optimization process. To evaluate CGMCL’s multimodal fusion performance, we tested various combinations of loss functions. These include the binary loss for single graph module fusion ($\mathcal{L}_{I}$ and $\mathcal{L}_{C}$), the addition of contrastive loss ($\mathcal{L}_{contrastive}$), and the comprehensive weighted loss CGMCL ($\mathcal{L}_{CGMCL}$). The experimental results in Fig. 4 show that $\mathcal{L}_{CGMCL}$ outperforms the combination of $\mathcal{L}_{I}$ + $\mathcal{L}_{C}$ by 2.1% in AUC and 2.2% in ACC for the MA vs. Abnormal classification task. Similarly, for the Normal vs. Abnormal task, $\mathcal{L}_{CGMCL}$ shows improvements of 1.1% in both AUC and ACC. These results indicate that the similarity between the two modalities for these subtypes is nearing the model’s convergence range. More importantly, the ability to distinguish early-stage PD is critical. In the Normal vs. MA classification task, $\mathcal{L}_{CGMCL}$ surpasses the next best-performing unweighted model $\mathcal{L}_{CGMCL}$ by 5.1% in AUC and 4.3% in ACC. This substantial improvement highlights CGMCL’s superior capability in capturing latent patterns characteristic of early-stage PD within the fusion space.